MIXED CYCLIC CONSTRAINTS ON CONFORMATIONAL FLEXIBILITY IN β/γ-PEPTIDES: CONFOR-MATION SPECIFIC IR AND UV SPECTROSCOPY

\underline{JOSEPH R. GORD}, PATRICK S. WALSH, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette. IN 47907; BRIAN F. FISHER and SAMUEL H. GELLMAN, Department of Chemistry, University of Wisconsin-Madison, Madison, WI 53706.

In order to further understand the intramolecular forces governing secondary structure formation in peptides and to provide benchmarks for the computational community, conformation-specific spectroscopy techniques have been applied to several model systems provided by Dr. Sam Gellman’s research group at the University of Wisconsin-Madison. In the present work, two model β/γ-peptides, Ac-β_{ACP C}-γ_{ACHC}-NHBz and Ac-γ_{ACHC}-β_{ACP C}-NHBz have been investigated using single and double resonance ultraviolet and infrared spectroscopy to elucidate their intrinsic folding propensities. The β-peptide is constrained by a five-membered ring spanning the β^{3}-β^{2} positions (β_{ACP C}) and the γ-peptide is constrained by a six-membered ring spanning the γ^{4}-γ^{3} positions with an additional ethyl group at γ^{2} (γ_{ACHC}). Resonant twophoton ionization (R2PI) spectra from 37250 to 37750 cm^{−1} were obtained and subsequently interrogated using UV-UV hole-burning to reveal the presence of three conformations for Ac-β_{ACP C}-γ_{ACHC}-NHBz, and a single conformation for Ac-γ_{ACHC}-β_{ACP C}-NHBz. Resonant ion-dip infrared (RIDIR) spectra were obtained in the NH stretch region from 3300 to 3500 cm^{−1} and in both the amide I and II regions from 1400 to 1800 cm^{−1}. These spectra were compared to computational predictions given by DFT calculations using the M05-2X functional with a 6-31G+(d) basis set revealing two slightly varied iterations of a bifurcated C-8/13 double ring structure for Ac-β_{ACP C}-γ_{ACHC}-NHBz and one bifurcated C-9/13 double ring structure for Ac-γ_{ACHC}-β_{ACP C}-NHBz. The appearance of C-13 rings was also seen in solution phase studies.^{a} This work is a complement to studies performed on pure γ-peptides and α/γ-peptides.

aL. Guo, A. M. Almeida, W. Zhang, A. G. Reidenbach, S. H. Choi, I.. A. Guzei, and S. H. Gellman J. Am. Chem. Soc. 2010, 132, 7868-7869